# Fast entangling gates on fluxoniums via parametric modulation of plasmon interaction

### v1

@@ I Introduction
For fluxonium qubits (the lowest five states labeled by $\{|0\rangle,|1\rangle,|2\rangle,|3\rangle,|4\rangle\}$), which are typically biased at the half-flux-quantum sweet spot to achieve high coherence times Nguyen2019 ; Somoroff2023 ; Wang2025 , the small electric dipole moment of the qubit transition $|0\rangle\rightarrow|1\rangle$ complicates its use in quantum operations beyond single-qubit gates, particularly within large-scale systems. However, among accessible higher-energy transitions, three plasmon transitions, namely $|0\rangle\rightarrow|3\rangle$, $|1\rangle\rightarrow|2\rangle$, and $|1\rangle\rightarrow|4\rangle$, possess transmon-like dipole moments, making them particularly relevant for implementing various quantum operations. Specifically, these transitions can facilitate strong dispersive coupling with readout resonators, enabling fast, high-fidelity qubit readout Stefanski2024 ; Bothara2025 . They can also, in principle, be leveraged to realize microwave-activated controlled-phase gates by facilitating transitions between computational states and non-computational states (involving the excitation of these plasmon modes Nesterov2018 or coupler modes Simakov2023 ). Apparently, both types of operations could inevitably cause leakage out of the computational subspace, presenting a significant challenge for quantum error correction Miao2023 . Addressing this issue requires the incorporation of leakage removal operations, typically involving engineered interactions between fluxonium plasmons and a dissipative environment such as the readout resonator Wang2024a . It is therefore evident that the practical utilization of plasmon modes in large-scale fluxonium-based quantum processors necessitates a flexible and scalable control strategy for engineering plasmon interactions Zhao2025 ; Zhao2025b .
@@
Here we introduce a control strategy that employs parametric modulation of the plasmon interaction to implement fast entangling gates on fluxonium qubits, alternative to the commonly used microwave-based approaches with static plasmon-plasmon couplings Nesterov2018 ; Ficheux2021 ; Xiong2022 ; Simakov2023 ; Ding2023 ; Rosenfeld2024 ; Xiong2025 ; Zhao2025 ; Singh2025 . Within the fluxonium qubit architecture supporting tunable plasmon interactions Zhao2025 , fluxoniums are coupled via a tunable coupler, see Fig. 1(a). The two-qubit gate can be implemented by applying parametric driving to the coupler at the sum frequency of the selective plasmon modes of the two fluxoniums (e.g., $|1\rangle\rightarrow|2\rangle$ for each fluxonium). This drive activates a bSWAP-type interaction Bertet2006 ; Niskanen2007 ; Roth2017 ; Poletto2012 ; Kapit2015 ; Zhao2017 ; Nesterov2021 (e.g., $|11\rangle\leftrightarrow|22\rangle$) between the selective plasmon modes, enabling the transition between computational states (e.g, $|11\rangle$) and the non-computational states (e.g, $|22\rangle$), see Fig. 1(b). Similar to the microwave-based approaches Nesterov2018 ; Simakov2023 , this transition enables the accumulation of a non-trivial conditional phase on the computational state (e.g, $|11\rangle$) and thus facilities the implementation of controlled-phase gates.
@@
Despite these advantages, the proposed strategy shares a significant challenge with conventional microwave-based methods, i.e., the short coherence times of plasmon modes (or coupler modes). Furthermore, since our approach involves doubly excited plasmon states, this coherence limitation is expected to be more severe compared to conventional schemes, which typically utilize only singly excited plasmon states. However, unlike parametric-activated gates in transmon systems that often exceed 100 ns in gate lengths McKay2016 ; Han2020 ; Li2022 ; Ganzhorn2020 , the present strategy enables sub-100ns CZ gates on fluxoniums with intrinsic errors below $10^{-4}$. We therefore expect that, even with coherence times on the order of $\sim 10\ \mu{\rm s}$, which aligns with typical values in current devices Ficheux2021 ; Ding2023 , a gate error approaching $10^{-3}$ remains achievable. Moreover, since the decoherence mechanisms of plasmon modes are similar to those of transmon qubits Ficheux2021 ; Nguyen2019 , one can anticipate no fundamental barrier to achieving coherence times approaching $100\ \mu{\rm s}$ in future implementations. Such an improvement could push gate errors toward the $10^{-4}$ level. Given all these considerations, we expect this approach may constitute a promising control strategy for the development of large-scale, high-performance fluxonium-based quantum processors.
@@
Figure 1: (a) A two-dimensional (2D) square qubit lattice comprising fluxoniums (circles) coupled via couplers (squares). The inset depicts the fluxonium architecture featuring tunable plasmon interactions, where fluxoniums are coupled via a frequency-tunable transmon coupler. (b) The energy levels of the unit cell comprising two coupled fluxonium qubits, with emphasis on the computational subspace spanned by $\{|000\rangle,|001\rangle,|100\rangle,|101\rangle\}$ (shaded region) and the fluxonium’s plasmon mode $|1\rangle\rightarrow|2\rangle$. The full system state is labeled as $|Q_{0},C,Q_{1}\rangle$. Solid orange arrows indicate direct plasmon-coupler couplings, while dashed orange arrows represent coupler-mediated plasmon-plasmon interactions. In addition to blue sideband transitions between each fluxonium and the coupler (red arrows), parametric modulation of the coupler can also activate a bSWAP-type interaction ($|101\rangle\rightarrow|202\rangle$, black arrow) between the plasmon modes of the two fluxoniums.
@@ IV CZ gate implementations based on the bSWAP interaction
Figure 7: Gate errors (a, c) and leakage (b, d) of the parametric CZ gates as functions of gate length. (a, b) and (c, d) correspond to the dynamic flux-bias configuration and the static bias configuration, respectively, see Fig. 2(b). Solid and dashed lines represent results for drive pulse ramp times of 5 ns and 10 ns, respectively. Gray dashed lines in (a) and (b) indicate the calculated incoherent gate errors due to relaxation and dephasing of the noncomputational state $|22\rangle$, assuming typical coherence times of $T_{1}^{22}=T_{2}^{22}=5\,\mu\rm s$.
@@ IV.1 Gate error
To characterize the intrinsic gate performance, we consider the metric of state-average gate fidelity Pedersen2007 . For each gate length and the coupler bias setting, optimal drive parameters, including the parametric drive frequency and amplitude, can be obtained by minimizing both leakage Wood2018 and conditional phase error Zhao2025 ; Jiang2025 (see Appendix C for detail). As shown in Figs. 7(a) and 7(b), the resulting gate error and leakage are shown as functions of gate length for both operational configurations. These results reveal a clear speed-fidelity trade-off: longer gate durations enable lower gate errors rates, and the gate performance is mainly limited by leakage into noncomputational states.
@@
Moreover, as indicated by the gray lines in Fig. 7, we estimate the contribution to gate error from the dominant decoherence channel, i.e., relaxation and dephasing of the noncomputational state $|22\rangle$ Zhao2025 ; Abad2023 (see Appendix for details), which is particularly relevant in practical implementations. This estimate assumes typical coherence times on the order of $\sim 10\ \mu{\rm s}$ for the plasmon mode $|1\rangle\rightarrow|2\rangle$, consistent with values reported in current devices Ficheux2021 ; Ding2023 . These results demonstrate that fast entangling gates could be realized with an error approaching $10^{-3}$, which is comparable to state-of-the-art experimental demonstrations Zhang2024 ; Lin2025 .
@@ Acknowledgements.
Peng Zhao would like to thank Zhuang Ma for insightful discussions. This work is supported by the National Natural Science Foundation of China (Grants No.12204050 and No.12275090) and the Guangdong Provincial Quantum Science Strategic Initiative (Grant No. GDZX2203001). Peng Xu is supported by the National Natural Science Foundation of China (Grants No.12105146 and No.12175104) and the Program of State Key Laboratory of Quantum Optics Technologies and Devices (No.KF202505).
@@ Appendix C The implementation of the Parametric-activated CZ Gate
where $U$ represents the truncated actual evolution operator within the computational subspace and $U_{\rm cz}$ corresponds to the ideal CZ gate. Here the actual evolution operator $U$ is obtained through numerical simulation of the gate dynamics, where each fluxonium is truncated to its five lowest energy levels and the transmon coupler is modeled as an anharmonic oscillator, see Eq. (14).
To assess incoherent errors in realistic devices, we note that in high-coherence fluxonium systems, as demonstrated in Ref. Ding2023 , the dominated incoherence gate error should be from relaxation and dephasing of non-computational gate states. We therefore focus specifically on the gate transition $|11\rangle\leftrightarrow|22\rangle$. In accordance with Refs. Abad2023 ; Zhao2025 , the incoherence gate error from the relaxation and dephasing (under white noise approximation) are
$\epsilon=1-F=\frac{3}{32}\frac{t_{g}}{T_{1}^{22}}+\frac{13}{80}\frac{t_{g}}{T_{\phi,{\rm white}}^{22}},$ (27)
where $T_{1}^{22}$ and $T_{\phi,{\rm white}}^{22}$ denote the relaxation time and the dephasing (white noise) time of the non-computational level $|22\rangle$.

### v2

@@ I Introduction
For fluxonium qubits (the lowest five states labeled by $\{|0\rangle,|1\rangle,|2\rangle,|3\rangle,|4\rangle\}$), which are typically biased at the half-flux-quantum sweet spot to achieve high coherence times Nguyen2019 ; Somoroff2023 ; Wang2025 , the small electric dipole moment of the qubit transition $|0\rangle\rightarrow|1\rangle$ complicates its use in quantum operations beyond single-qubit gates Moskalenko2021 ; Moskalenko2022 , particularly within large-scale systems. However, among accessible higher-energy transitions, three plasmon transitions, namely $|0\rangle\rightarrow|3\rangle$, $|1\rangle\rightarrow|2\rangle$, and $|1\rangle\rightarrow|4\rangle$, possess transmon-like dipole moments, making them particularly relevant for implementing various quantum operations. Specifically, these transitions can facilitate strong dispersive coupling with readout resonators, enabling fast, high-fidelity qubit readout Stefanski2024 ; Bothara2025 . They can also, in principle, be leveraged to realize microwave-activated controlled-phase gates by facilitating transitions between computational states and non-computational states (involving the excitation of these plasmon modes Nesterov2018 or coupler modes Simakov2023 ). Apparently, both types of operations could inevitably cause leakage out of the computational subspace, presenting a significant challenge for quantum error correction Miao2023 . Addressing this issue requires the incorporation of leakage removal operations, typically involving engineered interactions between fluxonium plasmons and a dissipative environment such as the readout resonator Wang2024a . It is therefore evident that the practical utilization of plasmon modes in large-scale fluxonium-based quantum processors necessitates a flexible and scalable control strategy for engineering plasmon interactions Zhao2025 ; Zhao2025b .
@@
Here we introduce a control strategy that employs parametric modulation of the plasmon interaction to implement fast entangling gates on fluxonium qubits, alternative to the commonly used microwave-based approaches with static plasmon-plasmon couplings Nesterov2018 ; Ficheux2021 ; Xiong2022 ; Simakov2023 ; Ding2023 ; Rosenfeld2024 ; Xiong2025 ; Zhao2025 ; Singh2025 . Within the fluxonium qubit architecture supporting tunable plasmon interactions Zhao2025 , fluxoniums are coupled via a tunable coupler, see Fig. 1(a). The two-qubit gate can be implemented by applying parametric driving to the coupler at the sum frequency of the selective plasmon modes of the two fluxoniums (e.g., $|1\rangle\rightarrow|2\rangle$ for each fluxonium). This drive activates a bSWAP-type interaction Bertet2006 ; Niskanen2007 ; Roth2017 ; Poletto2012 ; Kapit2015 ; Zhao2017 ; Nesterov2021 (e.g., $|11\rangle\leftrightarrow|22\rangle$) between the selective plasmon modes, enabling the transition between computational states (e.g, $|11\rangle$) and the non-computational states (e.g, $|22\rangle$), see Fig. 1(b). Similar to the microwave-based approaches Nesterov2018 ; Simakov2023 , this transition enables the accumulation of a non-trivial conditional phase on the computational state (e.g, $|11\rangle$) and thus facilities the implementation of controlled-phase gates. We note that, similar to the schemes studied in Refs. Moskalenko2021 ; Moskalenko2022 ; Didier2018 , employing parametric modulation on the fluxonium qubits could serve as an alternative approach for activating such bSWAP-type transitions.
@@
Despite these advantages, the proposed strategy shares a significant challenge with conventional microwave-based methods, i.e., the short coherence times of plasmon modes (or coupler modes). Furthermore, since our approach involves doubly excited plasmon states, this coherence limitation is expected to be more severe compared to conventional schemes, which typically utilize only singly excited plasmon states. However, unlike parametric gates in transmon systems, which often exceed 100 ns in gate lengths (see, e.g., Refs. McKay2016 ; Han2020 ; Li2022 ; Ganzhorn2020 ; a notable exception is the fast parametric scheme in Ref. Kubo2023 with gate times below 25 ns), the present strategy enables sub-100ns CZ gates on fluxoniums with intrinsic errors below $10^{-4}$. We therefore expect that, even with coherence times on the order of $\sim 10\ \mu{\rm s}$, which aligns with typical values in current devices Ficheux2021 ; Ding2023 , a gate error approaching $10^{-3}$ remains achievable. Moreover, since the decoherence mechanisms of plasmon modes are similar to those of transmon qubits Ficheux2021 ; Nguyen2019 , one can anticipate no fundamental barrier to achieving coherence times approaching $100\ \mu{\rm s}$ in future implementations. Such an improvement could push gate errors toward the $10^{-4}$ level. Given all these considerations, we expect this approach may constitute a promising control strategy for the development of large-scale, high-performance fluxonium-based quantum processors.
@@
Figure 1: (a) A two-dimensional (2D) square qubit lattice comprising fluxoniums (circles) coupled via couplers (squares). The inset depicts the fluxonium architecture featuring tunable plasmon interactions, where fluxoniums are coupled via a frequency-tunable transmon coupler. (b) The energy levels of the unit cell comprising two coupled fluxonium qubits (i.e., $Q_{0}$ and $Q_{1}$), with emphasis on the computational subspace spanned by $\{|000\rangle,|001\rangle,|100\rangle,|101\rangle\}$ (shaded region) and the fluxonium’s plasmon mode $|1\rangle\rightarrow|2\rangle$. The full system state is labeled as $|Q_{0},C,Q_{1}\rangle$. Solid orange arrows indicate direct plasmon-coupler couplings, while dashed orange arrows represent coupler-mediated plasmon-plasmon interactions. In addition to blue sideband transitions between each fluxonium and the coupler (red arrows), parametric modulation of the coupler can also activate a bSWAP-type interaction ($|101\rangle\rightarrow|202\rangle$, black arrow) between the plasmon modes of the two fluxoniums.
@@ IV CZ gate implementations based on the bSWAP interaction
Figure 7: (a, c) Intrinsic gate errors (excluding qubit and coupler relaxation and dephasing) and (b, d) leakage of the parametric CZ gates as functions of gate length. (a, b) and (c, d) correspond to the dynamic flux-bias configuration and the static bias configuration, respectively, see Fig. 2(b). Solid and dashed lines represent results for drive pulse ramp times of 5 ns and 10 ns, respectively.
@@ IV.1 Gate error
To characterize the intrinsic gate performance (excluding qubit and coupler relaxation and dephasing), we consider the metric of state-average gate fidelity Pedersen2007 , see Appendix C for detail. For each gate length and the coupler bias setting, optimal drive parameters, including the parametric drive frequency and amplitude, can be obtained by minimizing both leakage Wood2018 and conditional phase error Zhao2025 ; Jiang2025 (see Appendix C for detail). As shown in Figs. 7(a) and 7(b), the resulting gate error and leakage are shown as functions of gate length for both operational configurations. These results reveal a clear speed-fidelity trade-off: longer gate durations enable lower gate errors rates, and the gate performance is mainly limited by leakage into noncomputational states.
@@
Moreover, as shown in Fig. 8, we assess the impact of fluxonium and coupler relaxation and dephasing on the proposed gate implementation. We assume the computational states ($|0\rangle-|1\rangle$) of the fluxonium have relaxation and dephasing times of $T_{1}=1\,\rm ms$ and $T_{\phi}=0.2\,\rm ms$, consistent with recent experiments (Ref. Ding2023 ). For the fluxonium plasmon transitions (e.g., $|1\rangle-|2\rangle$, $|0\rangle-|3\rangle$, $|1\rangle-|4\rangle$) and the coupler transitions, we set $T_{1}=T_{2}$. Reflecting current technological capabilities, we further assume the fluxonium plasmon modes have a relaxation time half that of the transmon coupler. Using these parameters and solving the Lindblad master equation (see Appendix C for details) Wood2018 ; Zhao2022c , Figure 8 presents the total gate error, including all decoherence channels, as a function of the relaxation and dephasing times for 55-ns and 75-ns CZ gates in the two operational configurations (the drive pulse ramp time is 5 ns). For comparison, the intrinsic gate errors are also indicated for both configurations. It can be seen that, with typical coherence times on the order of $\sim 10\ \mu{\rm s}$ for the plasmon modes (consistent with values reported in current devices Ficheux2021 ; Ding2023 ), fast entangling gates can achieve errors approaching $10^{-3}$, which is comparable to state-of-the-art experimental demonstrations Zhang2024 ; Lin2025 .
To further identify the dominant error source, we also plot the errors under two simplified decoherence models: one including only relaxation and dephasing of the plasmon transition $|1\rangle-|2\rangle$ that participates in the activated bSWAP-type interaction ($|11\rangle\leftrightarrow|22\rangle$), and another including only decoherence of the computational states ($|0\rangle-|1\rangle$). As expected, for high-coherence fluxoniums, the leading gate error arises from relaxation and dephasing of the non-computational gate transition $|1\rangle\leftrightarrow|2\rangle$, consistent with earlier studies Ficheux2021 ; Ding2023 ; Abad2023 ; Zhao2025 .
Figure 8: Total gate error, gate error assuming only relaxation and dephasing of the plasmon transition $|1\rangle-|2\rangle$, and gate error assuming only relaxation and dephasing of qubit transition $|0\rangle-|1\rangle$ as function of the relaxation and dephasing times. For easy reference, the intrinsic gate errors are also indicated. Here, we assume $T_{1}=1\,\rm ms$ and $T_{\phi}=0.2\,\rm ms$ for the fluxonium computational states $|0\rangle-|1\rangle$, and set $T_{1}=T_{2}$ for fluxonium plasmon transitions ($|1\rangle-|2\rangle$, $|0\rangle-|3\rangle$, $|1\rangle-|4\rangle$) and the coupler. Additionally, the relaxation time of fluxonium plasmon modes is taken as half that of the transmon coupler. (a) and (b) correspond to the 55-ns CZ gate in dynamic flux-bias configuration and the 75-ns CZ gate in static bias configuration, respectively. Here, the drive pulse ramp time is 5 ns.
@@ IV.2 Leakage error analysis
IV.3 Sensitivity analysis for gate parameter fluctuations
Figure 10: Intrinsic gate error for 60-ns and 80-ns CZ gates in the two operational configurations versus drive-frequency detuning $\delta$ and normalized drive amplitude $\delta_{\Phi}/\delta_{\Phi 0}$ relative to their optimal values. We note that, at the optimal control points, the intrinsic errors are $6.9\times 10^{-6}$ and $1.9\times 10^{-4}$, respectively, as shown in Figs. 7(a) and 7(c).
Figure 11: Intrinsic gate error for 60-ns and 80-ns CZ gates in the two operational configurations versus qubit-bias offset $\Delta_{\Phi,Q}$ (applied to $Q_{0}$) and coupler-bias offset $\Delta_{\Phi,C}$ relative to their optimal values. We note that at the optimal control points, the intrinsic errors are $6.9\times 10^{-6}$ and $1.9\times 10^{-4}$, respectively, as shown in Figs. 7(a) and 7(c).
Fluctuations or drifts in system parameters (including those of both qubits and control electronics) can degrade gate performance (see, e.g., Ref. Ficheux2021 ; Ganzhorn2020 ). To evaluate their effect on intrinsic gate error, we analyze the sensitivity of the gate implementations to instabilities in various control parameters.
As an illustration, Figures 10 and 11 present results for 60-ns and 80-ns CZ gates in the two operational configurations (using a 5-ns drive-pulse ramp time). Note that, at the optimal control points, the intrinsic gate errors at the optimal control points are $6.9\times 10^{-6}$ and $1.9\times 10^{-4}$, respectively, as shown in Figs. 7(a) and 7(c). Figure 10 shows the influence of drive-parameter (frequency and amplitude) fluctuations on the gates, while Fig. 11 displays the effect of bias-parameter (flux biases on fluxonium $Q_{0}$ and the transmon coupler) fluctuations. As expected, fluctuations in system parameters do affect gate performance. At the same time, the results in these figures demonstrate that coherent gate errors below 0.001 are achievable with current experimental technology Ficheux2021 .
IV.4 Spectator-induced gate error
Figure 12: Spectator-induced gate errors for the CZ gate applied to $Q_{0}$ and $Q_{1}$ as a function of the residual coupling between $Q_{0}$ and spectator fluxonium $Q_{2}$ (controlled via the bias on the $Q_{0}-Q_{2}$ coupler; see Fig. 1(a)). Vertical dashed lines mark the bias points that suppress the residual coupling (see also in Fig. 2(a)) for suppressing the residuals. (a) 55-ns CZ gates in the dynamic flux-bias configuration with the spectator $Q_{2}$ prepared in different qubit states. (b) 75-ns CZ gates in the static flux-bias configuration with the spectator $Q_{2}$ prepared in different qubit states.
Here, to evaluate the extensibility of the proposed gate implementations, Figure 12 shows the intrinsic gate error while accounting for the residual coupling between $Q_{0}$ and spectator fluxonium $Q_{3}$, which is controlled by the coupler flux bias for the $Q_{0}-Q_{2}$ coupler (see Fig. 1(a)), for both operational configurations.
Following typical experimental protocols Krinner2020 ; Cai2021 , the CZ gate between $Q_{0}$ and $Q_{1}$ is first tuned with spectator fluxonium $Q_{2}$ in $|0\rangle$ and then characterized for both $|0\rangle$ and $|1\rangle$ states. As expected, stronger $Q_{0}-Q_{2}$ coupling generally increases the gate error for both spectator states, with a more pronounced degradation when the spectator is in $|1\rangle$. Significantly, when the qubit-spectator coupling is suppressed, the gate fidelity approaches that of an isolated two-qubit system, and the gate performance exhibits little dependence on the spectator state. Conversely, strong coupling can lead to frequency collisions and conditional frequency shifts (see Fig. 2(a) and Ref. Zhao2025 ), which substantially degrade the gate and introduce spectator-dependent correlated errors. These results, consistent with the analysis in Ref. Zhao2025 , underscore the key role of tunable plasmon interactions in mitigating spectator-induced errors.
@@ Acknowledgements.
Peng Zhao would like to thank Zhuang Ma for insightful discussions. This work is supported by the National Natural Science Foundation of China (Grants No.12204050 and No.92576110) and the Guangdong Provincial Quantum Science Strategic Initiative (Grant No. GDZX2203001). Peng Xu is supported by the National Natural Science Foundation of China (Grants No.12105146, No.12175104, and No.92565111) and the Program of State Key Laboratory of Quantum Optics Technologies and Devices (No.KF202505).
@@ Appendix C The implementation of the Parametric-activated CZ Gate
where $U$ represents the truncated actual evolution operator within the computational subspace and $U_{\rm cz}$ corresponds to the ideal CZ gate. Here the actual evolution operator $U$ is obtained through numerical simulation of the gate dynamics, where each fluxonium is truncated to its five lowest energy levels and the transmon coupler is described in a Fock basis Koch2007 with a dimension of 21.
We now assess the influence of fluxonium and coupler relaxation and dephasing on gate performance. Because the fidelity metric defined in Eq. (26) is not applicable under decoherence, we adopt the average gate fidelity defined as Wood2018 ; Zhao2022c
$\begin{aligned} F=\frac{4\,F_{p}+1-L_{1}}{5},\end{aligned}$ (27)
where $L_{1}$ denotes the leakage of the gate operation Wood2018 and $F_{p}$ is the process fidelity of the implemented CZ gate. Both quantities are obtained by numerically solving the Lindblad master equation Wood2018 ; Zhao2022c . Note that unlike the calculation of intrinsic gate errors, the transmon coupler is here modeled as a three-level anharmonic oscillator (see Ref. Zhao2025b ) to lower the computational overhead of solving the Lindblad equation, while still capturing the essential physics of the coupled fluxonium system (the same modeling approach is also used in the analysis of spectator-induced gate errors). This difference in modeling leads to distinct intrinsic gate errors for CZ gates of the same duration, i.e., the 55-ns and 75-ns CZ gates in the two operational configurations, as illustrated in Figs. 7 and 8.